# Origen: Zero-Shot 3D Orientation Grounding in Text-to-Image Generation

### v1

@@ 4 Experiments
Orientation Zero-1-to-3 [46] C3DW [11] Origen (Ours) Orientation Zero-1-to-3 [46] C3DW [11] Origen (Ours)
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3: Qualitative comparisons on MS-COCO-Single benchmark (Sec. 4.3). Compared to the existing orientation-to-image models [11, 46], Origen generates the most realistic images, which also best align with the grounding conditions, indicated by the overlapped arrow in each image.
Figure 4: Qualitative comparisons on MS-COCO-NView benchmark (Sec. 4.4). Origen generates high-fidelity images with the best grounding accuracy. Note that C3DW [11] does not support back-view generation, as it only supports the control of the front $180^{\circ}$ of azimuth angles.
Orientation Origen (Ours) Orientation Origen (Ours) Orientation Origen (Ours) Orientation Origen (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative results on the MS-COCO-Multi benchmark (Sec. 4.5) Origen achieves accurate grounding while generalizing across multiple objects.
@@ 4.1 Datasets
To the best of our knowledge, no existing benchmark has been proposed to evaluate 3D orientation grounding in text-to-image generation, aside from user studies conducted by Cheng et al. [11]. To address this, we introduce three benchmarks based on the MS-COCO dataset [42], that consist of diverse text prompts, image and bounding boxes.
MS-COCO-Single.
For a comparison with previous orientation-grounding methods [11, 46] that can condition on only a single object, we construct the MS-COCO-Single dataset. From MS-COCO validation set [42], we filter out (1) object classes for which orientation cannot be clearly defined (e.g., objects with front-back symmetry) and (2) image captions lacking explicit object references. Since the current orientation-to-image generation model [11] is only capable of controlling the front 180° range of azimuths, we further filter the samples to only include those within this range. This procedure yields 252 text-image pairs covering 25 distinct object classes. Using the provided bounding boxes, we cropped the foreground objects and fed them into OrientAnything [67], an orientation estimation model, to obtain pseudo-GT grounding orientations. Building upon this dataset, MS-COCO-Single was constructed by mix-matching the image captions and grounding orientations, ultimately forming a dataset consisting of 25 object classes, each with 40 samples, totaling 1K samples. Please refer Tab. A in Appendix to check object classes we used.
MS-COCO-NView.
One may consider achieving orientation grounding through text prompts for specific views. To further compare our method with such cases, we construct an extended dataset, MS-COCO-NView, based on the 252 text-image pairs from MS-COCO-Single. Specifically, by generating four grounding orientations – front, left, right, back – for each sample, we enable direct comparison with existing text-to-image models by appending these directional cues to prompts (e.g., adding facing left to the end of the prompt). In our experiments, we defined ground-truth (GT) orientations on a horizontal plane with four azimuth angles: $0^{\circ}$ (front), $90^{\circ}$ (left), $180^{\circ}$ (back), and $270^{\circ}$ (right). We consider two cases: (1) 3-View configuration that covers the front, left, and right views within a 180° frontal range (to further compare with C3DW [11]), (2) 4-View configuration that additionally includes a back view.
MS-COCO-Multi.
For more qualitative results, we construct MS-COCO-Multi following an approach to that in MS-COCO-Single. Since our base dataset [42] lacks samples composed solely of objects with a clear orientation, we mix-match object classes and grounding orientations from the 252 text-image pairs in MS-COCO-Single, forming a dataset consisting of 371 samples, each containing a varying number of objects. We annotated prompts by concatenating individual object phrases (e.g., “a cat, and a dog.").
@@ 4.2 Evaluation Metrics
We measure orientation grounding accuracy using two metrics: 1) the absolute error on azimuth angles [^3] We perform comparisons only on azimuth angle, as existing methods [11, 46] do not support the control over polar and rotation angles. Note that our results for all azimuth, polar, and rotation angles are provided in Appendix E. between the predicted and grounding object orientations, and 2) Acc.@22.5°, the angular accuracy within a tolerance of ±22.5°, following Wang et al. [67]. For evaluation, we use OrientAnything [67] to predict the 3D orientation from the generated images. Since OrientAnything’s predictions may not be perfect, we also conduct a user study in Sec. 4.6 to validate the results. Along with grounding accuracy, we also evaluate text-to-image alignment using CLIP Score [21] and VQA-Score [43], as well as human preference using PickScore [31].
4.3 Results on MS-COCO-Single
@@
Baselines. We compare Origen with two types of baselines: (1) orientation-to-image generation methods [11, 46] and (2) training-free guided generation methods [74, 16]. For orientation-to-image generation methods, we include Continuous 3D Words (C3DW) [11] and Zero-1-to-3 [46], following the baselines used in the most recent work in this field [11]. For training-free guided generation methods, we consider FreeDoM [74], which updates the intermediate samples at each step of the multi-step sampling process based on the expected reward computed on the foreseen clean samples, and ReNO [16], which optimizes the initial latent through vanilla gradient ascent using one-step models. To ensure fair comparisons, we keep the number of function evaluations (NFEs) consistent (NFEs = 50) across different guided generation methods. For example, in FreeDoM, we set the rewind iterations to 5 and inference steps to 10 to match 50 NFEs.
Results. In Tab. 1(a), we show our quantitative comparison results on the MS-COCO-Single benchmark. Origen significantly outperforms all the baselines in orientation alignment, showing comparable performance in text-to-image alignment. Our qualitative comparisons are also shown in Fig. 3, demonstrating that Origen generates high-quality images that align with the grounding orientation conditions and input text prompts. Note that C3DW [11] is trained on synthetic data (i.e., multi-view renderings of a 3D object) to learn orientation-to-image generation. Thus, it has limited generalizability to real-world images and the output images lack realism. Zero-1-to-3 [46] is also trained on single-object images but without backgrounds, requiring additional background image composition (also used in the evaluation of C3DW [11]) that may introduce unnatrual artifacts. The existing methods on guided generation methods also achieve suboptimal results compared to Origen. Notably, FreeDoM [74] achieves the worst results among the guided generation methods. This is because controlling the object orientation requires modifying the low-frequency structures of the image – which are known to be determined at early stages of sampling [12], as discussed in Sec. 2. ReNO [16] also achieves suboptimal results compared to ours, as it performs latent optimization based on vanilla gradient ascent which is prone to local optima (as discussed in Sec. 3.1). Overall, our method achieves the best results both with and without time rescaling, demonstrating its ability to effectively maximize rewards while avoiding over-optimization.
@@
Table 1: Quantitative comparisons on 3D orientation grounded image generation. Best and second-best results are highlighted in bold and underlined, respectively.
(a) Comparisons on MS-COCO-Single (Sec. 4.3). Origen denotes ours without reward-adaptive time rescaling, where Origen represents our full method.
Id Model Orientation Alignment Text Alignment
Acc.@22.5° $\uparrow$ Abs. Err. $\downarrow$ CLIP $\uparrow$ VQA $\uparrow$ PickScore $\uparrow$
@@
2 Zero-1-to-3 [46] 0.499 59.03 0.272 0.663 0.213
3 FreeDoM [74] 0.741 20.90 0.259 0.728 0.225
4 ReNO [16] 0.796 20.56 0.247 0.663 0.212
5 Origen (Ours) 0.854 18.28 0.265 0.732 0.224
6 Origen (Ours) 0.871 17.41 0.265 0.735 0.224
(b) Comparisons on MS-COCO-NView (Sec. 4.4).
Id Model 3-view Alignment 4-view Alignment
Acc.@22.5° $\uparrow$ Abs. Err. $\downarrow$ Acc.@22.5° $\uparrow$ Abs. Err. $\downarrow$
1 SD-Turbo [73] 0.257 75.09 0.244 78.47
2 SDXL-Turbo [57] 0.189 78.44 0.196 81.88
3 FLUX-Schnell [34] 0.312 75.04 0.424 60.26
4 C3DW [11] 0.504 53.53 - -
5 Zero-1-to-3 [46] 0.366 68.70 0.321 75.10
6 Origen (Ours) 0.824 20.99 0.866 17.45
(c) Comparisons on MS-COCO-Multi (Sec. 4.5).
Model Orientation Alignment Text alignment
Acc.@22.5° $\uparrow$ Abs. Err. $\downarrow$ CLIP $\uparrow$ VQA $\uparrow$ PickScore $\uparrow$
Origen (Ours) 0.634 34.27 0.281 0.764 0.219
4.4 Results on MS-COCO-NView
Baselines.
We also compare Origen on the MS-COCO-NView benchmark with the text-to-image (T2I) generation models, where their orientation condition is provided via the input text prompt. As baselines, we consider several one-step T2I generative models: SD-Turbo [73], SDXL-Turbo [57], and FLUX-Schnell [34]. In particular, we appended a phrase that specifies the object orientation at the end of each caption in MS-COCO-NView dataset. For more comprehensive comparisons, we also included the orientation-to-image baseline models (C3DW [11] and Zero-1-to-3 [46]) considered in Sec. 4.3 in this experiment as well.
Results. As shown in Tab. 1(b), Origen significantly outperforms all baseline models in orientation alignment. Note that, although FLUX-Schnell [34] achieves the highest alignment among the vanilla T2I models, Origen surpasses it by more than 2.5 times in the 3-view alignment setting (82.4% vs. 31.2%) and more than 2 times in the 4-view alignment setting (86.6% vs. 42.4%). This substantial margin highlights the inherent ambiguity and lack of precise control in vanilla T2I approaches, as orientation information embedded within textual descriptions is less explicit and reliable compared to direct orientation guidance. We further demonstrate the advantage of Origen through qualitative comparisons in Fig. 4. Vanilla T2I models frequently fail to adhere strictly to the desired orientation, even with directional phrases in text prompts. In contrast, Origen consistently generates images accurately aligned with the specified orientations. Please refer to Appendix F  that shows Origen also achieves comparable text alignment results compared to the vanilla T2I models.
4.5 Results on MS-COCO-Multi
We additionally show the results of Origen on the MS-COCO-Multi benchmark. Since no baseline is capable of multi-object orientation grounding, we exclusively assess our method. Our quantitative results are shown in Tab. 1(c) and the qualitative examples are shown in Fig. 5. These results further demonstrate that our approach is seamlessly generalizable to multiple objects by simply averaging the orientation grounding reward across multiple objects.
@@ 4.6 User Study
Our previous quantitative evaluations were performed by comparing the grounding orientations and orientations estimated from the generated images using OrientAnything [67]. While its orientation estimation performance is highly robust (as seen in all of our qualitative results), we additionally conduct a user study to further validate the effectiveness of our method based on human evaluation performed by 100 participants on Amazon Mechanical Turk [9]. Each participant was presented with the grounding orientation, the input prompt, and the images generated by (1) Zero-1-to-3 [46], (2) C3DW [11], and (3) Origen. Then, they were asked to select the image that best matches both the grounding orientations and the input text prompt – directly following the user study settings in C3DW [11]. In Tab. 2, Origen was preferred by 58.18% of the participants, clearly outperforming the baselines. For more details of this user study, refer to Appendix G.
@@ 5 Conclusion
We presented Origen, the first 3D orientation grounding method for text-to-image generation across multiple objects and diverse categories. To enable test-time guidance with a pretrained discriminative model, we proposed a novel sampling approach based on Langevin dynamics, which adds just one line of code to gradient ascent while theoretically guaranteeing a balance between orientation conditioning and image realism. Additionally, we introduced a reward-adaptive time rescaling method to accelerate convergence. Our experiments demonstrated significantly superior performance compared to previous fine-tuning-based methods, along with capabilities that were previously infeasible: conditioning on multiple objects, elevations, rotations, and 360 azimuths.
@@
A Analysis on Norm-Based Regularization
@@
$=\mathbf{x}_{t}+\eta_{1}\nabla\hat{\mathcal{R}}(\mathbf{x}_{t})$
$\quad+\eta_{2}\nabla\Big{[}(d-1)\log\|\mathbf{x}_{t}\|^{2}-\frac{1}{2}\|\mathbf{x}_{t}\|^{2}\Big{]}$
@@
$=\mathbf{x}_{t}+\nabla\Phi(\mathbf{x}_{t}).$ (11)
where $\Phi(\mathbf{x})=\eta_{1}\hat{\mathcal{R}}(\mathbf{x})+\eta_{2}K(\mathbf{x})$.
This optimization process is an Euler discretization (with $\delta t=1$) of the ODE
@@
However, as illustrated in Fig. 2 , this deterministic gradient ascent does not directly prevent the iterates from deviating substantially from the original latent distribution. Also, the process may collapse to local maxima–even ones where the reward is not sufficiently high [30]. Our empricial observations indicate that this approach is ineffective for our application, presumably because the reward function defined over the latent exhibits many local maxima, causing the deterministic ascent to converge to suboptimal solutions.
B Proofs
@@
then $q^{*}(\mathbf{x})$ is the unique stationary distribution. Using Eq. 2 , we obtain
@@
which matches the expression presented in Eq. 2 . ∎
C Analysis on Reward-Adaptive Time-Rescaled SDE
In this section, we analyze the effect of a position-dependent step size on reward-guided Langevin dynamics. We first illustrate why a naive time-rescaling approach fails to preserve the desired stationary distribution, and how to fix it with an additional correction term, eventually leading to Eq. 8 . Our derivation follows the approach of Leroy et al. [39].

### v2

@@ 4 Experiments
Orientation Zero-1-to-3 [17] C3DW [18] FreeDoM [35] ReNO [22] Origen (Ours)
Figure 3: Qualitative comparisons on ORIBENCH-Single benchmark (Sec. 4.5). Compared to the existing orientation-to-image models [18, 17], Origen generates the most realistic images, which also best align with the grounding conditions in the leftmost column.
Orientation DPS [32] MPGD [33] FreeDoM [35] ReNO [22] Origen (Ours)
Figure 4: Qualitative comparisons on ORIBENCH-Multi benchmark (Sec. 4.5). Compared to the guided-generation methods [32, 33, 35, 22], Origen generates the most realistic images, which also best align with the grounding conditions in the leftmost column.
@@ 4.1 Datasets
To the best of our knowledge, no existing benchmark has been proposed to evaluate 3D orientation grounding in text-to-image generation, aside from user studies conducted by Cheng et al.  [18]. To address this, we introduce ORIBENCH based on the MS-COCO dataset [23], that consist of diverse text prompts, image, and bounding boxes.
ORIBENCH-Single.
For a comparison with previous orientation-grounding methods [18, 17] that can condition on only a single object, we construct the ORIBENCH-Single dataset. From MS-COCO validation set [23], we filter out (1) object classes for which orientation cannot be clearly defined (e.g., objects with front-back symmetry) and (2) image captions lacking explicit object references. Since the current orientation-to-image generation model [18] is only capable of controlling the front 180° range of azimuths, we further filter the samples to only include those within this range. This procedure yields 252 text-image pairs covering 25 distinct object classes. Using the provided bounding boxes, we cropped the foreground objects and fed them into OrientAnything [20], an orientation estimation model, to obtain pseudo-GT grounding orientations. Building upon this dataset, ORIBENCH-Single was constructed by mix-matching the image captions and grounding orientations, ultimately forming a dataset consisting of 25 object classes, each with 40 samples, totaling 1K samples (See Appendix E to check object classes we used).
ORIBENCH-Multi.
For the comparison of a complex scenario with the grounding orientation of multiple objects, we construct ORIBENCH-Multi following an approach to that in ORIBENCH-Single. Since our base dataset [23] lacks samples composed solely of objects with a clear orientation, we mix-match object classes and grounding orientations from the 252 non-overlapping text-image pairs in ORIBENCH-Single, forming a dataset consisting of 371 samples, each containing a varying number of objects. We annotated prompts by concatenating individual object phrases (e.g., “a cat, and a dog.”).
@@ 4.2 Evaluation Metrics
We compare Origen’s performance from two perspectives: (1) Orientation Alignment and (2) Text-to-Image Alignment. We measure orientation alignment using two metrics: 1) Acc.@$22.5^{\circ}$, the angular accuracy within a tolerance of ±$\pm 22.5^{\circ}$ and 2) the absolute error on azimuth angles [^1] We perform comparisons only on azimuth angle, as existing methods [18, 17] do not support the control over polar and rotation angles. Note that our results for all azimuth, polar, and rotation angles are provided in Appendix G.1. between the predicted and grounding object orientations, following Wang et al. [20]. For evaluation, we use OrientAnything [20] to predict the 3D orientation from the generated images. Since its predictions may not be perfect, we also conduct a user study in Sec. 4.6 to validate the results. Along with grounding accuracy, we also evaluate text-to-image alignment using CLIP Score [66], VQA-Score [67], and PickScore [68].
4.3 Baselines
@@
We compare Origen with three types of baselines: (1) training-based orientation-to-image generation methods [18, 17] (2) training-free guided generation methods [35, 22] for the multi-step generation, and (3) training-free guided generation methods for the one-step generation. For (1), we include Continuous 3D Words (C3DW) [18] and Zero-1-to-3 [17], following the baselines used in the most recent work in this field [18]. For (2), we consider DPS [32], MPGD [33], and FreeDoM [35], which update the intermediate samples at each step of the multi-step sampling process based on the expected reward computed on the estimated clean image. Finally, for (3), we compare Origen with ReNO [22], which optimizes the initial latent through vanilla gradient ascent using one-step models. Furthermore, to evaluate the effectiveness of our reward adaptive time rescaling, we perform additional comparison with our method variant without this component, denoted as Origen.
4.4 Implementation Details
We use FLUX-Schnell [21] as the one-step generative model for both Origen and ReNO [22], while all multi-step guided generation baselines (DPS [32], MPGD [33], and FreeDoM [35]) are implemented using FLUX-Dev [21] as the multi-step generative model. Except for the training-based methods (C3DW [18] and Zero-1-to-3 [17]), we match the number of function evaluations (NFEs—defined as the number of iterations in our method and the denoising steps in multi-step generative models) to 50 for a fair comparison across all methods. All experiments were conducted on a single NVIDIA 48GB VRAM A6000 GPU. Further implementation details are provided in Appendix F.
@@ 4.4 Implementation Details
Table 1: Quantitative comparisons on 3D orientation grounded image generation. Best and second-best results are highlighted in bold and underlined, respectively. Origen denotes ours without reward-adaptive time rescaling.
ORIBENCH-Single ORIBENCH-Multi
Id Model Orientation Alignment Text-to-Image Alignment Orientation Alignment Text-to-Image Alignment
(1) Fine-tuned Orientation-to-Image Model
1 Zero-1-to-3 [17] 0.499 59.03 0.272 0.663 0.213
@@
(2) Guided-Generation Methods with Multi-Step Model
3 DPS [32] 0.664 28.16 0.246 0.662 0.221 0.487 38.28 0.285 0.742 0.231
4 MPGD [33] 0.689 27.62 0.246 0.661 0.222 0.532 35.03 0.283 0.739 0.232
5 FreeDoM [35] 0.741 20.90 0.259 0.728 0.225 0.591 31.77 0.279 0.783 0.227
(3) Guided-Generation Methods with One-Step Model
6 ReNO [22] 0.796 20.56 0.247 0.663 0.212 0.478 45.95 0.277 0.756 0.223
7 Origen (Ours) 0.854 18.28 0.265 0.732 0.224 0.679 29.7 0.287 0.804 0.225
8 Origen (Ours) 0.871 17.41 0.265 0.735 0.224 0.692 28.4 0.287 0.807 0.225
4.5 Results
ORIBENCH-Single.
In the left part of Tab. 1, we show our quantitative comparison results on the ORIBENCH-Single benchmark. Origen significantly outperforms all the baselines in orientation alignment, showing comparable performance in text-to-image alignment. Our qualitative comparisons are also shown in Fig. 3, demonstrating that Origen generates high-quality images that align with the orientation conditions and input text prompts. Note that C3DW [18] is trained on synthetic data (i.e., multi-view renderings of a 3D object) to learn orientation-to-image generation. Thus, it has limited generalizability to real-world images and the output images lack realism. Zero-1-to-3 [17] is also trained on single-object images but without backgrounds, requiring additional background image composition (also used in the evaluation of C3DW [18]) that may introduce unnatrual artifacts. The existing methods on guided generation methods also achieve suboptimal results compared to Origen. In particular, all training-free multi-step guidance methods, including DPS [32], MPGD [33], and FreeDoM [35], perform poorly in orientation grounding. This is because object orientation control relies on modifying low-frequency image structures, which are primarily formed during early sampling stages, where multi-step diffusion models produce noisy and blurry outputs [69] (as discussed in Sec. 3.1). ReNO [22] also achieves suboptimal results compared to ours, as it performs latent optimization based on vanilla gradient ascent which is prone to local optima (as discussed in Sec. 3.1). Overall, our method achieves the best results both with and without time rescaling, demonstrating its ability to effectively maximize rewards while avoiding over-optimization.
ORIBENCH-Multi.
We additionally show the quantitative comparison results on the ORIBENCH-Multi benchmark. Since no fine-tuned methods ((1) in Tab. 1) are capable of multi-object orientation grounding, we assess our method only with guided generation methods ((2), (3) in Tab. 1). Our quantitative results are presented in the right of Tab. 1, and qualitative examples are provided in Fig. 4. As shown, Origen consistently outperforms all guided generation baselines under the same reward function. These results demonstrate that our object-agnostic reward design, combined with the proposed method, enables robust and generalizable orientation grounding even in complex multi-object scenarios.
Additional Results.
We provide additional results for orientation grounding task (1) under more diverse orientations, and (2) for four primitive views (front, left, right, back) to enable comparisons with text-to-image models, where the orientation condition is included as an input text prompt. Please refer to Appendix  G.
@@ 4.6 User Study
Our previous quantitative evaluations were performed by comparing the grounding orientations and orientations estimated from the generated images using OrientAnything [20]. While its orientation estimation performance is highly robust (as seen in all of our qualitative results), we additionally conduct a user study to further validate the effectiveness of our method based on human evaluation performed by 100 participants on Amazon Mechanical Turk [70]. Each participant was presented with the grounding orientation, the input prompt, and the images generated by (1) Zero-1-to-3 [17], (2) C3DW [18], and (3) Origen. Since Origen outperformed all guided-generation methods with same reward function design in quantitative comparison, they were excluded in the user study. Then, participants were asked to select the image that best matches both the grounding orientations and the input text prompt – directly following the user study settings in C3DW [18]. In Tab. 2, Origen was preferred by 58.18% of the participants, clearly outperforming the baselines. For more details of this user study, refer to Appendix H.
4.7 Computation Time
In our experiments, we set the number of iterations in our method to match the denoising steps in multi-step generative models (e.g., 50), resulting in comparable computation time—approximately 52.7 seconds per image. A detailed comparison is provided in Appendix I. Notably, for cases achieving angular accuracy within a tolerance of ±$\pm 22.5^{\circ}$ (Tab. 3), the average number of iterations to reach this threshold was 12.8, corresponding to an average of 13.5 seconds. Our approach can also serve as a post-refinement method for existing models, potentially further reducing computation time.
@@ 5 Conclusion
We presented Origen, the first 3D orientation grounding method for text-to-image generation across multiple objects and categories. To enable test-time guidance with a pretrained discriminator, we introduced a Langevin-based sampling approach with reward-adaptive time rescaling for faster convergence. Experiments showed that our method outperforms fine-tuning-based baselines and uniquely supports conditioning on multiple open-vocabulary objects.
Limitations and Societal Impact.
Since our method relies on a pretrained discriminator, its performance is inherently bounded by the quality of that discriminator—though in our case, it still achieved state-of-the-art results by a significant margin. As a generative AI technique, our method may be misused to produce inappropriate or harmful content, underscoring the importance of responsible development and deployment.
@@
Appendix A Analysis on Norm-Based Regularization
@@
$=\mathbf{x}_{t}+\eta_{1}\nabla\hat{\mathcal{R}}(\mathbf{x}_{t})+\eta_{2}\nabla\Big{[}(d-1)\log\|\mathbf{x}_{t}\|^{2}-\frac{1}{2}\|\mathbf{x}_{t}\|^{2}\Big{]}$
@@
$=\mathbf{x}_{t}+\nabla\Phi(\mathbf{x}_{t})$ (11)
where $\Phi(\mathbf{x})=\eta_{1}\hat{\mathcal{R}}(\mathbf{x})+\eta_{2}K(\mathbf{x})$. This optimization process is an Euler discretization (with $\delta t=1$) of the ODE
@@
However, as illustrated in Fig. 2 in the main paper, this deterministic gradient ascent does not directly prevent the iterates from deviating substantially from the original latent distribution. Also, the process may collapse to local maxima–even ones where the reward is not sufficiently high [71]. Our empricial observations indicate that this approach is ineffective for our application, presumably because the reward function defined over the latent exhibits many local maxima, causing the deterministic ascent to converge to suboptimal solutions.
Appendix B Proofs
@@
then $q^{*}(\mathbf{x})$ is the unique stationary distribution. Using Eq. 2 in the main paper, we obtain
@@
which matches the expression presented in Eq. 2 in the main paper. ∎
Appendix C Analysis on Reward-Adaptive Time-Rescaled SDE
In this section, we analyze the effect of a position-dependent step size on reward-guided Langevin dynamics. We first illustrate why a naive time-rescaling approach fails to preserve the desired stationary distribution, and how to fix it with an additional correction term, eventually leading to Eq. 8 in the main paper. Our derivation follows the approach of Leroy et al. [65].